# SeNeDiF-OOD: Semantic Nested Dichotomy Fusion for Out-of-Distribution Detection Methodology in Open-World Classification. A Case Study on Monument Style Classification

### v2

@@ 5.3.2 Global results
ID class: known-style monuments.
@@ 5.4 Images filtered by layers
This section provides a qualitative, image-based view of how the cascade behaves in practice. Here, we examine concrete inputs that either slip through a layer or that progressively survive all filters. The goal is twofold: first, to distill characteristic visual patterns that explain the few errors that remain after each gate, and second, to illustrate how the hierarchical fusion cascade progressively narrow the operational design domain from generic imagery to known-style monuments. These observations will be revisited in the discussion and used to motivate targeted data curation and training refinements.
@@ 5.4.1 Representative misclassifications per layer
(h) Non monument with monument like frontage
@@ 6 Conclusions
In this work, we have introduced SeNeDiF-OOD, a novel methodoloy that operationalizes OOD detection and open-world recognition by integrating semantic nested dichotomies into a robust decision fusion architecture. The approach is motivated by the inherent open-world setting of applications such as MonuMAI, where user-submitted inputs frequently extend beyond the intended scope of the model. Our contribution has been twofold. At the theoretical level, we have formalized a taxonomy of OOD categories and established a general framework for applying nested dichotomies to image-based OOD detection. At the practical level, we have implemented this methodology within the MonuMAI pipeline, designing and training dedicated filtering layers that progressively refine the decision boundaries.
@@ Acknowledgements
This research results from the Strategic Project IAFER-Cib (C074/23), as a result of the collaboration agreement signed between the National Institute of Cybersecurity (INCIBE) and the University of Granada. This initiative is carried out within the framework of the Recovery, Transformation and Resilience Plan funds, financed by the European Union (Next Generation).

### v3

@@ 5.3.2 Global results
in-distribution class: known-style monuments.
@@ 5.4 Images filtered by layers
This section provides a qualitative, image-based view of how the cascade behaves in practice. Here, we examine concrete inputs that either slip through a layer or that progressively survive all filters. The goal is twofold: first, to distill characteristic visual patterns that explain the few errors that remain after each gate, and second, to illustrate how the hierarchical fusion cascade progressively narrows the operational design domain from generic imagery to known-style monuments. These observations will be revisited in the discussion and used to motivate targeted data curation and training refinements.
@@ 5.4.1 Representative misclassifications per layer
(h) Non-monument with monument-like frontage
@@ 6 Conclusions
In this work, we have introduced SeNeDiF-OOD, a novel methodology that operationalizes OOD detection and open-world recognition by integrating semantic nested dichotomies into a robust decision fusion architecture. The approach is motivated by the inherent open-world setting of applications such as MonuMAI, where user-submitted inputs frequently extend beyond the intended scope of the model. Our contribution has been twofold. At the theoretical level, we have formalized a taxonomy of OOD categories and established a general framework for applying nested dichotomies to image-based OOD detection. At the practical level, we have implemented this methodology within the MonuMAI pipeline, designing and training dedicated filtering layers that progressively refine the decision boundaries.
@@ Acknowledgements
This research was supported by the Strategic Project IAFER-Cib [C074/23] through a collaboration agreement signed between the National Institute of Cybersecurity (INCIBE) and the University of Granada. This initiative is carried out within the framework of the Recovery, Transformation and Resilience Plan funds, financed by the European Union (Next Generation EU).